The excavation for a swimming pool of 5m by 10m, is 8ft. deep.
1. How many cubic meters of soil is to be excavated?
2. What is the lateral area of the pool?
3. How much is the total excavation cost at Php 500 per cubic
meter of earth removed?
1
Expert's answer
2020-10-08T16:35:22-0400

The swimming pool is a cuboid with length 5m,width 10m and height 8ft

Converting feet into meters ;1ft=0.3048m therefore 8ft=(8*0.3048)m=2.4384m

1.cubic meters of soil to be excavated :

"V=l*b*h" where "l=length"

"b=width, h=height"

"V=5m*10m*2.4384m=121.92m^3"

2.lateral area =2(l+b)h

"=2(5m+10m)2.4384m=73.152m^2"

3.total cost :

500Php for every "m^3" excavated

Therefore total cost "=121.92m^3*500Php\/m^3=60960Php"